# Complementarity between quantum entanglement, geometrical and dynamical appearances in N spin-1/2 system under all-range Ising model

### v2

@@ I Introduction
The main purpose of this work is to exploit the geometrical structures caracterizing the quantum phase space of $N$ interacting spin-$1/2$ under all-range Ising model to explore some physical properties, such as the geometric phase, the curvature of quantum phase space, the evolution speed, distance geodesic between quantum states, solving quantum brachistochrone problem, and their interplay with quantum entanglement. It is noteworthy that the ideas explored in this paper were primarily inspired by the findings obtained by Krynytskyi and Kuzmak in Ref.[27]. As a matter of fact, by performing the system dynamics, we derive the Fubini-Study metric identifying the associated quantum state space. Moreover, examining the Gaussian curvature (G-curvature) in the framework of the Gauss-Bonnet theorem, we determine the topology and the structure of this space. Afterward, we explore the acquired geometric and topological phases and tackle the quantum brachistochrone issue. Finally, we give a detailed explanation of the geometrical and dynamical characteristics of two interacting spin-$1/2$ under the Ising model in connection to quantum entanglement.
@@ II.1 Physical model and unitary quantum evolution
$\mathtt{S}_{i}^{z}\left|{{\mathsf{m}_{i}}}\right\rangle=\hbar{\mathsf{m}_{i}}\left|{{\mathsf{m}_{i}}}\right\rangle,$ (2)
@@
$\left[{\mathtt{S}_{i}^{\alpha},\mathtt{S}_{j}^{\beta}}\right]=i{\delta_{ij}}\sum\limits_{\gamma=x,y,z}\epsilon^{\alpha\beta\gamma}{\mathtt{S}_{k}^{\gamma}},$ (3)
@@
$\left|\Psi(t)\right\rangle=\sum_{p=0}^{N}\cos^{N-p}\left(\frac{\theta}{2}\right)\sin^{p}\left(\frac{\theta}{2}\right)\exp\left\{-i\left[\frac{\xi(t)}{4}(N-2p)^{2}-p\varphi\right]\right\}\sum_{i_{1}<i_{2}<\ldots<i_{p}=1}^{\mathcal{N}}\sigma_{i_{1}}^{x}\sigma_{i_{2}}^{x}\ldots\sigma_{i_{p}}^{x}|\uparrow\uparrow\ldots\uparrow\rangle,$ (7)
with $\xi(t)=\mathtt{J}t$. In this way, we come at demonstrating the evolved state of a collection of $N$ qubits (i.e., $N$ spin1/212-1/2- 1 / 2) that is explicitly dependent on three parameters, namely the spherical angles $(\theta,\varphi)$ and the time $t$. It is also intriguing to note that the state (7) fulfills the periodic requirement $\left|\Psi(\xi)\right\rangle=\left|\Psi(\xi+2\pi)\right\rangle$ along the temporal parameter, implying that $\xi\in[0,2\pi]$. Accordingly, one can predict that the dynamics of the system occurs on a closed three-dimensional manifold.
@@ II.2 Geometry and topology of the resulting state manifold
After evolving the $N$ spin-$1/2$ system by means of temps-evolution operator and determining the evolved state (7), we will now discover the geometry and topology associated with the relevant quantum state space, which includes all possible states that the system may reach during evolution. For this task, we must compute the Fubini-Study metric, which defines the infinitesimal distance $d\mathtt{S}$ between two adjoining pure quantum states $|\Psi(\zeta^{\mu})\rangle$ and $|\Psi(\zeta^{\mu}+d\zeta^{\mu})\rangle$, having the following form [54, 55]
@@
where $\zeta^{\mu}$ are the physical parameters $\theta,\varphi$ and $\xi$ (i.e., representing the dynamical degrees of freedom of the considered system) specifying the evolved state (7) and $\mathrm{g}_{\mu\nu}$ denote the components of this metric tensor given by
@@
with $\left|\Psi_{\mu,\nu}\right\rangle=\frac{\partial}{\partial\zeta^{\mu,\nu}}|\Psi\rangle$. Using the definition (8), one finds, after a simple numerical calculation, the explicit version of the Fubini-Study metric as follows
@@
To further characterize this state space, we are going to determine its topology. For this aim, we begin by assessing the corresponding G-curvature, which can be defined in terms of the relevant metric tensor (12) in the form [56]
@@
Remark that the state space curvature (15) is mainly affected by the initial parameters $\theta$ and $N$, while it is independent of $\xi$ containing the temporal evolution, meaning that the state space curvature is independent of the system dynamics. Furthermore, the G-curvature (15) verifies the following periodic requirement $\mathrm{K}(\theta)=\mathrm{K}(\theta+\pi)$. This is consistent with our findings because the resulting quantum phase space (12) is a closed two-dimensional manifold.
@@
where the geometric invariants $d\mathrm{S}$, $\mathrm{k}_{\mathrm{g}}$ and $d{l}$ denote, respectively, the surface element, geodesic curvature and line element. Additionally, the first and second terms on the left-hand side of the equation (16) represent, respectively, the bulk and border contributions to the Euler characteristic identifying the state space topology. Furthermore, the Gauss-Bonnet theorem (16) can be established in terms of the state space geometry (12) in the form
@@
here $\Lambda$ standing for the Euler border integral including the conical flaws contribution. After a simple calculation, the Gauss-Bonnet theorem (17) reads as
@@
As a further note, the solid angle of a revolution cone having the angle at the peak $2\theta$ is defined by $\Omega=2\pi(1-\cos\theta)$ where the second term of this definition represents the partial solid angle sketched out by the system around the cone apex during evolution. Taking into account the closeness of the angular flaws at the two singular points, we can easily find
$2\pi\cos\theta\approx\frac{\mathtt{S}\left(2\pi\right)}{\mathbf{d}}=\frac{2\pi\sqrt{\mathrm{g}_{\xi\xi}}}{\sqrt{\mathrm{g}_{\theta\theta}}\theta},$ (20)
here $\mathtt{S}(2\pi)$ standing for the distance traveled by the system within the period of time $\textbf{t}=2\pi/\mathtt{J}$ around one of the two singular positions $(\theta=0$ or $\theta=\pi)$, while d corresponds to the distance between evolution of the system and the relevant singular position. As a result, the angular flaws are explicitly given by
@@
where we multiplied it by the factor 2 because we have two singular points. Putting the equation (21) into (18), one finds the Euler characteristic $\chi(\mathrm{M})=2$, showing that the quantum phase space associated with the $N$ spin-$1/2$ system has a spherical topology. The coming section will be devoted to a thorough analysis of the geometric phase that the $N$ spin-$1/2$ state (7) can accumulate when subjected to cyclic and arbitrary evolution processes on the underlying quantum state space (12).
@@ III Geometrical phases acquired by the $N$ spin-$1/2$ state
After investigating the geometry and topology of the $N$ spin-$1/2$ state space identified by the metric tensor (12). Let us now focus on the geometric phase that the evolving state (7) can acquire for both arbitrary and cyclic evolutionary processes.
@@ III.2 Geometrical phase under a cyclic evolution
here $|\tilde{\Psi}(t)\rangle$ denotes the Anandan-Aharonov section given in Ref. [36]. It is identified through the evolved state (7) as $|\tilde{\Psi}(t)\rangle=e^{-i{f}(t)}|\Psi(t)\rangle$ where ${f}(t)$ is any smooth function checking $f({T})-f(0)=\Phi_{\operatorname{tot}}$. On this view, the AA-geometric phase (31) rewrites
@@ IV.1 Speed of quantum evolution
This shows that the dynamics of the system occurs on a circle of radius $\sqrt{\mathrm{g}_{\theta\theta}}$. Therefore, the evolution speed of the $N$ spin-$1/2$ state (7) takes the form [9]
@@ IV.2 Resolution of the quantum brachistochrone problem
$\mathrm{t}_{\min}=\frac{\mathtt{S}_{\min}}{\mathrm{V}_{\max}}=\frac{t}{(N-1)}\sqrt{2N-3}.$ (44)
@@ V.2 Geometrical picture of the entanglement
In the same framework, we can also investigate the influence of entanglement on the G-curvature of the two-spin phase space (50). As a matter of fact, putting the equation (49) into (15), we derive the curvature according to the concurrence under the form
@@ V.3 Dynamical picture of the entanglement
Thereby, we arrive at expressing the Fubini-Study distance (i.e., dynamical observable), relating any two quantum states over the two-spin phase space (50), in terms of the entanglement level and the evolution time. This proves again the feasibility to investigate experimentally the dynamical properties, which will motivate their exploitation in the novel applications of quantum technology [63].
@@
By scrutinizing the two figures (7) and (8), we find that the Fubini-Study distance (61) exhibit, with respect to the entanglement degree, the same behavior as the evolution velocity (60), and thus the same conclusions can be achieved. Let us now solve the quantum brachistochrone issue for the two interacting spins based on their entanglement. To realize this, maximizing the evolution speed (60) with respect to the concurrence, the shortest time required to achieve a time-optimal evolution, over the relevant phase space (50), reads as
@@ VI Conclusion
To summarize, we investigated a physical system consisting of $N$ interacting spin-$1/2$ under all-range Ising model. We assumed that the starting state is the tensor product of $N$ eigenstates of the spin-$1/2$ projection operator on the positive direction denoted by the unit vector. After applying the time evolution propagator, the obtained evolved state is identified by three dynamical degrees of freedom, which are the spherical angles $(\theta,\varphi)$ and the evolution time $t$. We established the Fubini-Study metric and explored that the related quantum phase space is a closed two-dimensional manifold. Moreover, by examining the G-curvature in the framework of the Gauss-Bonnet theorem, we demonstrated that this state space has both a dumbbell-shape structure and a spherical topology. The geometric phase acquired by the $N$ spin-$1/2$ system is also studied within the arbitrary and cyclic evolution processes. This is achieved by evaluating the difference between the total and dynamic phases. We found that the geometric phase, in the arbitrary evolution case, varies nonlinearly with the time, reflecting its dynamic character. Geometrically, we showed that it is affected both by the evolution trajectory taken by the system and the associated state space geometry. On this view, we concluded that the geometric phase can be exploited to parameterize the possible evolution trajectories of the system. This result may be interesting for building robust quantum gates. Further, in the thermodynamic limit ($N\to\infty$), the total phase cancels out and therefore the geometric and dynamic phases coincide at any moment in the evolution process. This offers the opportunity of measuring the geometric phase experimentally. In the cyclic evolution case, we have calculated the AA-geometric phase and discovered that it is independent of the system dynamics. Otherwise, it depends only on the initial state choice (i.e., the initial parameters), signifying that it is influenced by the state space geometry and not by the evolution path followed by the system. Hence the cyclic evolution paths are not parameterizable by the AA-geometric phase. We have also derived the topological phase appearing naturally during cyclic evolutions. We found that it is proportional to the square of the particle number $N^{2}$, especially it takes fractional values for $N$ odd and multiples of $\pi$ for $N$ even. This signifies that the number of spins influences the topology of the state space. The evolution speed and the Fubini-Study distance separating the quantum states are also well examined. As a result, we resolved the quantum Brachistochrone problem for the $N$ spin-$1/2$ system by determining the shortest time (i.e., optimal time) required to perform an time-optimal evolution, and defined the underlying optimal state space. In this perspective, we discovered that for $N=2$ (i.e., two spin1/212-1/2- 1 / 2 system) the optimal and ordinary times coincide $(\mathrm{t}_{\min}=t)$, while for $N\geq 3$ (i.e., $N$ spin-$1/2$ system), the optimal time (44) is strictly lower than the ordinary time, and thus the time-optimal evolution is achievable. Besides, in the thermodynamic limit $(N\to\infty)$, the optimal time decreases to zero $(\mathrm{t}_{\min}\to 0)$. In this scheme, the optimal state circle coincides with a straight line since its radius becomes infinite. Thereby, we inferred that the particle number and the ordinary time are two physical magnitudes exploitable for performing the time-optimal evolutions in such integrable systems.
@@
On other hand, by restricting the whole system to a two-spin system, i.e., two interacting spin under the Ising model, we have studied the quantum entanglement via two approaches. The first approach is of geometric nature, in which we give the Fubini-Study metric in connection to the Wootters concurrence as a quantum correlation quantifier. This outcome may be interesting for the experimental handling of the geodesic distance between entangled states and also in the state space geometry adjustment. We proved that an increase in the entanglement degree between the two spins causes a decrease in the state space curvature until it reaches negative values, showing the quantum correlation effect in the compactification of the related state space. Additionally, the entanglement can be used to identify the quantum states over the space of states, for example, the states of maximum entanglement are housed in the regions of lowest curvature, whereas the separable states are housed in the regions of highest curvature. The geometric phase acquired by the two-spin system is sufficiently discussed in relation to the entanglement. We explored that the geometric phase exhibits two different behaviors with respect to the critical entanglement level $\boldsymbol{\mathscr{C}}_{\text{c}}$ $(\Phi_{\operatorname{g}}(\boldsymbol{\mathscr{C}}_{\text{c}})={\Phi_{\operatorname{g}}}_{\min})$: the first one is in the interval $[0,\boldsymbol{\mathscr{C}}_{\text{c}}]$, wherein the quantum correlations favors the loss of the geometric phase, while the second one is in the interval $[\boldsymbol{\mathscr{C}}_{\text{c}},1]$, wherein the quantum correlations favors the gain of the geometric phase. In the cyclic evolution scenario, the AA-geometric phase behaves similarly to the geometric phase in the first interval $[0,\boldsymbol{\mathscr{C}}_{\text{c}}]$. This highlights the significance of quantum entanglement for controlling the geometric phase evolvement in such spin systems. The second approach is of dynamic nature, we linked the evolution speed with the concurrence, we observed that the speed displays two different behaviors with respect to the critical entanglement level $\boldsymbol{\mathscr{C}}_{\text{c}}^{\prime}$ $(\mathrm{V}(\boldsymbol{\mathscr{C}}_{\text{c}}^{\prime})=\mathrm{V}_{\max})$: the first one is in the interval $[0,\boldsymbol{\mathscr{C}}_{\text{c}}^{\prime}]$, wherein the quantum correlations speed up the system evolution over the relevant phase space, while the second one is in the interval $[\boldsymbol{\mathscr{C}}_{\text{c}}^{\prime},1]$, wherein the quantum correlations slow down this evolution. Accordingly, we concluded that the system dynamics is controllable by its entanglement level. The same behavior is noticed for the Fubini-Study distance separating the entangled states. Finally, we solved the quantum barchistochrone problem based on the entanglement amount exchanged between the two spins. We inferred that the quantum entanglement and the ordinary time are two physical magnitudes exploitable to realize the time-optimal evolution in such a spin system. Thus, we were able to illustrate, to a significant extent, the connection between quantum entanglement and the geometric and dynamical characteristics characterizing the considered two-spin phase space.

### v3

@@ I Introduction
The main purpose of this work is to exploit the geometrical structures characterizing the quantum phase space of $N$ interacting spin-$1/2$ under all-range Ising model to explore some physical properties, such as the geometric phase, the curvature of quantum phase space, the evolution speed, distance geodesic between quantum states, solving quantum brachistochrone problem, and their interplay with quantum entanglement. It is noteworthy that the ideas explored in this paper were primarily inspired by the findings obtained by Krynytskyi and Kuzmak in Ref.[27]. As a matter of fact, by performing the system dynamics, we derive the Fubini-Study metric identifying the associated quantum state space. Moreover, examining the Gaussian curvature (G-curvature) in the framework of the Gauss-Bonnet theorem, we determine the topology and the structure of this space. Afterward, we explore the acquired geometric and topological phases and tackle the quantum brachistochrone issue. Finally, we give a detailed explanation of the geometrical and dynamical characteristics of two interacting spin-$1/2$ under the Ising model in connection to quantum entanglement.
@@ II.1 Physical model and unitary quantum evolution
$\mathtt{S}_{i}^{z}\left|{{\mathsf{m}_{i}}}\right\rangle={\mathsf{m}_{i}}\left|{{\mathsf{m}_{i}}}\right\rangle,$ (2)
@@
$\left[{\mathtt{S}_{i}^{\alpha},\mathtt{S}_{j}^{\beta}}\right]=i{\delta_{ij}}\sum\limits_{\gamma=x,y,z}\epsilon^{\alpha\beta\gamma}{\mathtt{S}_{i}^{\gamma}},$ (3)
@@
$\left|\Psi(t)\right\rangle=\sum_{p=0}^{N}\cos^{N-p}\left(\frac{\theta}{2}\right)\sin^{p}\left(\frac{\theta}{2}\right)\exp\left\{-i\left[\frac{\xi(t)}{4}(N-2p)^{2}-p\varphi\right]\right\}\sum_{i_{1}<i_{2}<\ldots<i_{p}=1}^{N}\sigma_{i_{1}}^{x}\sigma_{i_{2}}^{x}\ldots\sigma_{i_{p}}^{x}|\uparrow\uparrow\ldots\uparrow\rangle,$ (7)
where we define $\xi(t)=\mathtt{J}t$ as the dynamical parameter of the system. In this way, we come at demonstrating the evolved state of a collection of $N$ qubits (i.e., $N$ spin1/212-1/2- 1 / 2) that is explicitly dependent on three parameters, namely the spherical angles $(\theta,\varphi)$ and the dynamical parameter $\xi$. It is also intriguing to note that the state (7) fulfills the periodic requirement $\left|\Psi(\xi)\right\rangle=\left|\Psi(\xi+2\pi)\right\rangle$ along the temporal parameter, implying that $\xi\in[0,2\pi]$. Accordingly, one can predict that the dynamics of the system occurs on a closed three-dimensional manifold.
@@ II.2 Geometry and topology of the resulting state manifold
After evolving the $N$ spin-$1/2$ system by means of evolution operator and determining the evolved state (7), we will now discover the geometry and topology associated with the relevant quantum state space, which includes all possible states that the system may reach during evolution. For this task, we must compute the Fubini-Study metric, which defines the infinitesimal distance $d\mathtt{S}$ between two adjoining pure quantum states $|\Psi(\zeta^{\mu})\rangle$ and $|\Psi(\zeta^{\mu}+d\zeta^{\mu})\rangle$, having the following form [54, 55]
@@
where $\zeta^{\mu}$ are the physical parameters $\theta,\varphi$ and $\xi$ specifying the evolved state (7) and $\mathrm{g}_{\mu\nu}$ denote the components of this metric tensor given by [54]
@@
with $\left|\Psi_{\mu}\right\rangle=\frac{\partial}{\partial\zeta^{\mu}}|\Psi\rangle$. Using the definition (8), one finds, after a simple numerical calculation, the explicit version of the Fubini-Study metric as follows
@@
To further characterize this state space, we are going to determine its topology. For this aim, we begin by assessing the corresponding G-curvature, measuring the intrinsic curvature of the resulting quantum state manifold (12). It can be defined in terms of the relevant metric tensor (12) in the form [56]
@@
Note that the state space curvature (15) is mainly affected by the initial parameters $\theta$ and $N$, while it is independent of $\xi$ containing the temporal evolution, meaning that the state space curvature is independent of the system dynamics. Furthermore, the G-curvature (15) verifies the following periodic requirement $\mathrm{K}(\theta)=\mathrm{K}(\theta+\pi)$. This is consistent with our findings because the resulting quantum phase space (12) is a closed two-dimensional manifold.
@@
where the geometric invariants $d\mathrm{S}$, $\mathrm{k}_{\mathrm{g}}$ and $d{l}$ denote, respectively, the surface element, geodesic curvature, i.e., the curvature at the boundary of the state space (12), and line element. Additionally, the first and second terms on the left-hand side of the equation (16) represent, respectively, the bulk and border contributions to the Euler characteristic identifying the state space topology. Furthermore, the Gauss-Bonnet theorem (16) can be established in terms of the state space geometry (12) in the form
@@
here $\Lambda$ denotes the Euler border integral including the conical flaws contribution. After a simple calculation, the Gauss-Bonnet theorem (17) reads as
@@
It follows that
@@
where we multiplied it by the factor 2 because we have two singular points. The detailed derivation of the result (20) is provided in the appendix. Putting the equation (20) into (18), one finds the Euler characteristic $\chi(\mathrm{M})=2$, showing that the quantum phase space associated with the $N$ spin-$1/2$ system has a spherical topology. The coming section will be devoted to a thorough analysis of the geometric phase that the $N$ spin-$1/2$ state (7) can accumulate when subjected to cyclic and arbitrary evolution processes on the underlying quantum state space (12).
@@ III Geometrical phases acquired by the $N$ spin-$1/2$ state
After investigating the geometry and topology of the $N$ spin-$1/2$ state space identified by the metric tensor (12), let us now focus on the geometric phase that the evolving state (7) can acquire for both arbitrary and cyclic evolutionary processes.
@@ III.2 Geometrical phase under a cyclic evolution
here $|\tilde{\Psi}(t)\rangle$ denotes the Anandan-Aharonov section given in Ref. [36]. It is identified through the evolved state (7) as $|\tilde{\Psi}(t)\rangle=e^{-i{f}(t)}|\Psi(t)\rangle$ where ${f}(t)$ is any smooth function satisfying $f({T})-f(0)=\Phi_{\operatorname{tot}}$. On this view, the AA-geometric phase (30) rewrites
@@ IV.1 Speed of quantum evolution
This shows that the dynamics of the system occurs on a circle of radius $\sqrt{\mathrm{g}_{\xi\xi}}$. Therefore, the evolution speed of the $N$ spin-$1/2$ state (7) takes the form [9]
@@ IV.2 Resolution of the quantum brachistochrone problem
$\mathrm{t}_{\min}=\frac{\mathtt{S}_{\min}}{\mathrm{V}_{\max}}=\frac{t}{(N-1)}\sqrt{2N-3},\quad\text{ with}\,N>1.$ (44)
@@ V.2 Geometrical picture of the entanglement
In the same framework, we can also investigate the influence of entanglement on the G-curvature of the two-spin phase space (50). As a matter of fact, putting the equation (49) into (15), we derive the G-curvature according to the concurrence under the form
@@ V.3 Dynamical picture of the entanglement
Thereby, we arrive at expressing the Fubini-Study distance (i.e., dynamical observable), relating any two quantum states over the two-spin phase space (50), in terms of the entanglement level and the evolution time. This proves again the feasibility to investigate experimentally the dynamical properties, which will motivate their exploitation in the novel applications of quantum technology [63, 64].
@@
By scrutinizing the two figures (7) and (8), we find that the Fubini-Study distance (61) exhibits, with respect to the entanglement degree, the same behavior as the evolution velocity (60), and thus the same conclusions can be achieved. Let us now solve the quantum brachistochrone issue for the two interacting spins based on their entanglement. To realize this, maximizing the evolution speed (60) with respect to the concurrence, the shortest time required to achieve a time-optimal evolution, over the relevant phase space (50), reads as
@@ VI Conclusion
To summarize, we investigated a physical system consisting of $N$ interacting spin-$1/2$ under all-range Ising model. We assumed that the starting state is the tensor product of $N$ eigenstates of the spin-$1/2$ projection operator on the positive direction denoted by the unit vector. After applying the time evolution propagator, the obtained evolved state is identified by three dynamical degrees of freedom, which are the spherical angles $(\theta,\varphi)$ and the dynamical parameter $\xi$ (7). We established the Fubini-Study metric and explored that the related quantum phase space is a closed two-dimensional manifold (12). Moreover, by examining the G-curvature in the framework of the Gauss-Bonnet theorem, we demonstrated that this state space has both a dumbbell-shape structure and a spherical topology. The geometric phase acquired by the $N$ spin-$1/2$ system is also studied within the arbitrary and cyclic evolution processes (III.1). This is achieved by evaluating the difference between the total and dynamic phases. We found that the geometric phase, in the arbitrary evolution case, varies nonlinearly with the time, reflecting its dynamic character. Geometrically, we showed that it is affected both by the evolution trajectory taken by the system and the associated state space geometry. On this view, we concluded that the geometric phase can be exploited to parameterize the possible evolution trajectories of the system. This finding could be crucial for developing robust quantum gates and, consequently, for creating effective quantum circuits essential for quantum computation. Further, in the thermodynamic limit ($N\to\infty$), the total phase cancels out and therefore the geometric and dynamic phases coincide at any moment in the evolution process. This offers the opportunity of measuring the geometric phase experimentally. In the cyclic evolution case, we have calculated the AA-geometric phase and discovered that it is independent of the system dynamics (32). Otherwise, it depends only on the initial state choice (i.e., the initial parameters), signifying that it is influenced by the state space geometry and not by the evolution path followed by the system. Hence the cyclic evolution paths are not parameterizable by the AA-geometric phase. We have also derived the topological phase appearing naturally during cyclic evolutions. We found that it is proportional to the square of the particle number $N^{2}$, especially it takes fractional values for $N$ odd and multiples of $\pi$ for $N$ even. This signifies that the number of spins influences the topology of the state space. The evolution speed and the Fubini-Study distance separating the quantum states are also well examined (38). As a result, we resolved the quantum Brachistochrone problem for the $N$ spin-$1/2$ system by determining the shortest time (i.e., optimal time) required to perform an time-optimal evolution (44). Such evolutions play an essential role in advancing quantum circuits by reducing the effects of decoherence and ensuring the reliability of algorithms. In this perspective, we discovered that for $N=2$ (i.e., two spin1/212-1/2- 1 / 2 system) the optimal and ordinary times coincide $(\mathrm{t}_{\min}=t)$, while for $N\geq 3$ (i.e., $N$ spin-$1/2$ system), the optimal time (44) is strictly lower than the ordinary time, and thus the time-optimal evolution is achievable. Besides, in the thermodynamic limit $(N\to\infty)$, the optimal time decreases to zero $(\mathrm{t}_{\min}\to 0)$. In this scheme, the optimal state circle coincides with a straight line since its radius becomes infinite. Thereby, we deduced that the particle number and the ordinary time are two physical magnitudes exploitable for performing the time-optimal evolutions in such solvable systems.
@@
On other hand, by restricting the whole system to a two-spin system, i.e., two interacting spin under the Ising model, we have studied the quantum entanglement via two approaches. The first approach is of geometric nature, in which we give the Fubini-Study metric in connection to the Wootters concurrence as a quantum correlation quantifier (50). This outcome may be interesting for the experimental handling of the geodesic distance between entangled states and also in the state space geometry adjustment. We proved that an increase in the entanglement degree between the two spins causes a decrease in the state space curvature until it reaches negative values, showing the quantum correlation effect in the compactification of the related state space (see Fig. (4)). Additionally, the entanglement can be used to identify the quantum states over the space of states, for example, the states of maximum entanglement are housed in the regions of lowest curvature, whereas the separable states are housed in the regions of highest curvature. The geometric phase acquired by the two-spin system is sufficiently discussed in relation to the entanglement (57). We explored that the geometric phase (see Fig. (6)) exhibits two different behaviors with respect to the critical entanglement level $\boldsymbol{\mathscr{C}}_{\text{c}}$ $(\Phi_{\operatorname{g}}(\boldsymbol{\mathscr{C}}_{\text{c}})={\Phi_{\operatorname{g}}}_{\min})$: the first one is in the interval $[0,\boldsymbol{\mathscr{C}}_{\text{c}}]$, wherein the quantum correlations favors the loss of the geometric phase, while the second one is in the interval $[\boldsymbol{\mathscr{C}}_{\text{c}},1]$, wherein the quantum correlations favors the gain of the geometric phase. In the cyclic evolution scenario, the AA-geometric phase behaves similarly to the geometric phase in the first interval $[0,\boldsymbol{\mathscr{C}}_{\text{c}}]$. This highlights the significance of quantum entanglement for controlling the geometric phase evolvement in such spin systems. The second approach is of dynamic nature, we linked the evolution speed with the concurrence (60), we observed that the speed (see Fig. (7) displays two different behaviors with respect to the critical entanglement level $\boldsymbol{\mathscr{C}}_{\text{c}}^{\prime}$ $(\mathrm{V}(\boldsymbol{\mathscr{C}}_{\text{c}}^{\prime})=\mathrm{V}_{\max})$: the first one is in the interval $[0,\boldsymbol{\mathscr{C}}_{\text{c}}^{\prime}]$, wherein the quantum correlations speed up the system evolution over the relevant phase space, while the second one is in the interval $[\boldsymbol{\mathscr{C}}_{\text{c}}^{\prime},1]$, wherein the quantum correlations slow down this evolution. Accordingly, we concluded that the system dynamics can be controlled by its entanglement degree, which is a physical resource that can be measured experimentally. The same behavior is noticed for the Fubini-Study distance separating the entangled states (see Fig. (8)). Finally, we solved the quantum barchistochrone problem based on the entanglement amount exchanged between the two spins. We inferred that the quantum entanglement and the ordinary time are two physical magnitudes exploitable to realize the time-optimal evolution in such a spin system. Thus, we were able to illustrate, to a significant extent, the connection between quantum entanglement and the geometric and dynamical characteristics characterizing the considered two-spin phase space.
Appendix A
In this appendix we demonstrate the result (20), which determines the contribution of the angular defects in the Gauss-Bonnet theorem (17). The quantum state manifold defined by metric tensor (12) possesses a two singular points at $\theta=0$ and $\pi$. Near these points, the manifold exhibits an angular defects having the form of cones. Remark that if we rewrite the metric tensor for these areas, since these areas are close to the singular points, we can express the metric up to the second order in $\theta$ as in (19). Based on these considerations, we shall derive the equation that allows us to evaluate the angular defect near a singular point. The angular defect can be identified as a difference between the full angle $2\pi$ and the angle that the system traverses during one complete rotation (i.e., one period) around the singular point
$\Lambda=2\pi-\frac{\mathtt{S}({\xi_{\max}})}{d}$ (65)
where $\xi_{\max}=2\pi$, and $\mathtt{S}({\xi_{\max}})$ represents the distance that the system traverses during one complete rotation around the singular point and $d$ is the distance between the system path and the singular point. Now, to compute $\mathtt{S}(2\pi)/d$ we employ the expression (35) by taking $g_{\xi\xi}$ from the metric tensor (19). Thus, during the one complete rotation $\left(t=\xi_{\max}/\mathtt{J}\right)$ the system covers the following distance
$\mathtt{S}_{\xi_{\max}}=\sqrt{g_{\xi\xi}}\xi_{\max}=\frac{1}{2}(N-1)\sqrt{N}\theta\xi_{\max}.$ (66)
Taking $g_{\theta\theta}$ from the metric (19), we give $d$ as follows
$d=\sqrt{g_{\theta\theta}}\theta=\frac{1}{2}\sqrt{N}\theta.$ (67)
By introducing the equations (66) and (67) into (65), and considering that the manifold defined by the metric (12) possesses two angular defects, we find, after a simple calculation, the result (20).